# Maxime Dieudonné1, Guillaume Auzias1,†, Julien Lefèvre1,†

### v1

@@ Abstract
The shape of human brain is complex and highly variable, with interactions between brain size, cortical folding, and age well-documented in the literature. However, few studies have explored how global brain size influences geometric features of the cortical surface derived from anatomical MRI. In this work, we focus on sulcal depth, an imaging phenotype that has gained significant attention in both basic research and clinical applications. We make key contributions to the field by: 1) providing the first quantitative analysis of how brain size affects sulcal depth measurements; 2) introducing a novel, scale-invariant method for sulcal depth estimation based on an original formalization of the problem; 3) presenting a validation framework and sharing our code and benchmark data with the community; and 4) demonstrating the biological relevance of our new sulcal depth measure using a large sample of 1,987 subjects spanning the developmental period from 26 weeks post-conception to adulthood. Keywords. Magnetic resonance Imaging, cortical surface, sulcal depth
@@ 1 Introduction
Sulcal depth is an imaging phenotype that has attracted increasing attention across a wide range of studies, from fundamental research to clinical applications. Key applications include inter-individual cortical surface registration and atlas construction [Dale et al.,, Lyttelton et al.,, Robinson et al.,], detection of abnormal brain morphology related to clinical conditions [Auzias et al., b, Dierker et al.,], and exploration of the anatomo-functional relationships [Weiner and Golarai,]. Furthermore, sulcal depth plays a crucial role in studying individual variations in cortical folding at various spatial scales, from sulci [Duchesnay et al.,] to sulcal basins [Auzias et al., a, Im et al., a], and even more specific geometrical structures such as plis-de-passages [Bodin et al.,, Mangin et al., b] or gyral peaks [Zhang et al.,]. Despite its recognized importance for characterizing brain morphology, the sulcal depth is still underutilized compared to other widely used morphological descriptors, such as the gyrification index [Schaer et al.,, Rabiei et al.,], curvature-related characteristics [Pienaar et al.,], and sulcus-based morphometry [Mangin et al., a]. This limited use can be largely attributed to the inconsistent definitions of sulcal depth used in different studies, which complicates a unified understanding of the literature. Moreover, current methods for estimating sulcal depth often overlook the potential influence of variations in global brain size.
1.1 On the complex development of the cortical surface
The shape of the adult human brain cortex is usually described as a labyrinthine structure composed of ridges and valleys. This complex geometry is the result of the progressive folding that occurs mainly during the third trimester of gestation [Kostović et al.,]. The biological mechanisms underlying brain morphogenesis remain largely elusive [Llinares-Benadero and Borrell,, de Vareilles et al.,]. At the scale of millimeters that is accessible in vivo e.g. using MRI, the development of the morphology of the cortex is characterized by two intricate processes: global growth and folding. The causal relationship between variations in global brain size and cortical folding is thus still unknown.
@@ 1.2 On the magnitude of variations in global brain size
The magnitude of variations in global brain size has been extensively described in the literature. The recent large-scale study [Bethlehem et al.,] aggregated morphometry measures extracted from a very large population spanning the entire life, using image processing pipelines such as Freesurfer [Dale et al.,], or the dHCP pipeline [Makropoulos and Robinson,]. The authors applied advanced non-linear regression techniques to estimate the influence of the age on these morphometry measures. From these curves representative of a large population, we can observe a  4-fold increase in total brain volume between birth and adulthood. In a recent work investigating brain maturation before birth using MRI acquired in utero, [Studholme et al.,] reported a  12-fold increase between the $18^{th}$ week of gestation and birth. So regarding the human brain, variations related to brain maturation can reach an order of magnitude of  48-fold. Regarding inter-individual variations, we can also observe from [Bethlehem et al.,] a  3-fold variation in total brain volume ($[0.5,1.5]$ L) across healthy adults brains. On the side of cross-species comparisons, studies such as [Van Essen et al.,] reported  17-fold variation in total brain volume between mature macaques and mature humans. This brief literature overview is sufficient to remind that taking into account the variations in global brain size when comparing measures across brains is critical in most applications, either basic or clinically oriented.
@@ 1.3 Features estimated locally are affected by global brain size
It is also crucial to keep in mind that the relationship between global brain size and local measures of brain shape is not restricted to a linear interaction. Studies such as [Reardon et al.,] characterized these complex interactions and reported regionally specific patterns of non linear areal scaling in the human cortex across healthy adults. Such interactions between size and shape (often termed allometry) are very common in biological systems [Goriely,]. More specifically, systematic interactions between brain size, cortical folding and age have been extensively reported in the literature, both before and after birth [Rodriguez-Carranza et al.,, Yun et al., b, Garcia et al.,, Xu et al.,, Bozek et al.,]. Such interactions require particular attention in the context of developmental diseases that affect both brain size and cortical folding [Kapellou et al.,, Germanaud et al., a]
@@ 1.4 State-of-the-art of sulcal depth estimation techniques
The most intuitive strategy consists in defining the sulcal depth at each point of a cortical surface by computing the distance between the spatial location of the point of interest and the closest point on the ’convex hull’ of the brain. Already with this intuitive definition, two main ambiguities emerge when moving from the concept to its implementation. First, computing the ’convex hull’ that implicitly corresponds to the 0-level (or reference-level) for the depth measure is not trivial. Indeed, a mathematically correct definition of the convex hull of the cortical surface would not follow closely most concave regions such as the orbital frontal region and the internal temporal region, resulting in a local over estimation of the sulcal depth compared to e.g. the external, more convex regions. The available implementations thus rely instead on a proxy computed by applying a morphological closing of the volumetric segmentation mask of the cortex, which is efficient but question the potential influence between the size of the closing operator and the global size of the brain of interest. Second, there is no unique way to compute the distance to the convex hull and several approaches have been proposed, such as the Euclidean distance in 3D [Van Essen,], the geodesic distance along the cortical surface [Cachia et al.,], or an adaptive distance in 3D that follows the geometry of the surface [Yun et al., a]. We refer the readers interested in a detailed comparison across these types of approaches to [Yun et al., a] which provides both qualitative and quantitative comparisons.
A second strategy for estimating sulcal depth without considering the convex hull has been introduced in [Dale et al.,]. We refer to this method as SULC in the following. SULC is computed iteratively through an inflation process deforming an initial surface into a smooth, inflated one. Formally, the SULC estimation is obtained as:
@@
where $N_{i}(t)$ is the normal at $i$, $S_{i}(t)$ represents the infinitesimal deformation at vertex $i$ and time $t$, obtained through the gradient of an energy preserving local distances $d_{iv}=||e_{iv}||$ where $e_{iv}$ is the vector between vertex $i$ and one of its neighbours $v$
@@
A third strategy for estimating sulcal depth is based on two hypotheses relating sulcal depth and mean curvature: (1) the location of the extrema of mean curvature are a first approximation of the location of the extrema of sulcal depth and (2) the frequency ad amplitude of curvature and sulcal depth are closely related ([Luders et al.,], [Chung et al.,]).
Finally, a specific method denoted as Depth Potential Function (DPF) has been introduced in [Boucher et al.,] as a combination of the second and third strategies. The DPF has been defined as the solution $D$ of the following regularised screened Poisson equation:
@@
In addition to the specific limitations mentioned above, all the methods from the literature have been designed and evaluated on adult brains only. To the best of our knowledge, the potential influence of global brain size on the sulcal depth estimation obtained at each vertex of the cortical surface was not considered in any of these methods. The interactions between global brain size and cortical folding magnitude has been explored rarely, with the notable exception of the excellent work [Pienaar et al.,], which focuses on curvature but not sulcal depth. In application studies, global brain size is usually considered as a confounding factor to be controlled as much as possible in statistical models.There is thus a clear need to explicitly investigate more specifically the problem of the intricate variations in global brain size and folding.
1.5 The problem of the validation of sulcal depth estimation techniques
A key factor to explain the lack of a consensual definition of sulcal depth is the circularity between the concept of sulcal depth; and the need, for the validation of estimation techniques on real data, to define anatomical objects on the cortical surface that are not independent from sulcal depth, such as sulcus fundi, gyral crests or sulcal basins. Breaking this circularity would require a better understanding the biological mechanisms underlying brain folding. In absence of such precious knowledge, it is critical to limitate the circularity by stating as explicitly as possible the assumptions that support the definition of the geometrical objects used to assess the biological relevance of the estimated sulcal depth. The other possible path to advance in this interdisciplinary topic consists in improving the formalization of the problem of sulcal depth estimation.
In the present work, we contribute to this field of research by 1) providing the first quantitative characterisation of the influence of brain size on sulcal depth estimations; 2) introducing a new scale-invariant sulcal depth estimation method based on an original formalization of the problem; 3) proposing a validation framework and sharing with the community the benchmark data; 4) demonstrating the biological significance of our new sulcal depth measure using a large sample of 1,987 subjects, spanning the developmental period from 26 weeks post-conception to adulthood. The source code of this work is available at https://github.com/maximedieudonne/DPF-star, and the benchmark data is available at https://doi.org/10.5281/zenodo.14265228.
@@ 2.1 Scale controlled family of functions on surfaces
Note that the scale invariance from Definition 1 is not very relevant in applications to real brains. Indeed, as stated in the section 1.1 of the introduction, variations across any two brains in global brain size are systematically associated with variations in folding magnitude. When considering two cortical surfaces from different individuals, variations in brain size cannot be approximated by such simple scaling since brain volume and surface are related by an allometry. Even when considering the cortical surfaces from a single individual acquired at two different developmental stages, brain maturation impacts simultaneously both the scaling and the folding. As a consequence, in the vast majority of applications, the experimenter aims at comparing the functions computed on $M_{1}$ and $M_{2}$, knowing that $M_{2}\neq sM_{1}$.
@@ 2.3 Adaptation of the scale of interest through $\alpha$
Now let us focus on the setting of the only one parameter of DPF (Eq.3), and thus of $DPF^{*}$ : $\alpha$. As mentioned in [Boucher et al.,], decomposing the DPF into the eigenfunctions basis of Laplace-Beltrami operator (Fourier modes) allows to better interpret the influence of $\alpha$. The eigenpairs $(\lambda_{i},\Phi_{i})$ are solutions of
@@
This formal analysis allows us to interpret the $DPF^{\star}$ as a low-pass filter on the curvature, the level of which is controlled by $\alpha$. Decreasing $\alpha$ increases the influence of the low frequency components of the curvature compared to high frequency components, and vice versa. The low frequency components correspond to geometric information relative to the global shape of the brain, with large concave regions at the scale of a lobe (such as the prefrontal frontal and occipital regions). Reducing the high frequency components reduces the influence of local folds and noise within the sulci. We refer to [Germanaud et al., b] for further interpretation of $\lambda_{i}$ in terms of wavelengths of an eigenfunction $\Phi_{i}$.
@@ 3 Materials and Experiments
We evaluate our method using quantitative and qualitative comparison with the closest related approach from the literature, the method SULC that was described in Sc.1.4. In this work, we compute SULC using the deform-mesh process available in the bioMedIA mirtk pipeline  [^1] https://github.com/BioMedIA/MIRTK. We confirmed that the estimation is very close to the one from Freesurfer by computing the correlation between the two implementations for several brains and observed a correlation coefficient systematically superior to .99. We present four complementary experiments:
@@ 3.1 Data
1) The third release of the dHCP dataset (http://www.developingconnectome.org [Edwards and Rueckert,]) consists of 887 MRI sessions from 783 newborn babies covering the ages from 26 to 45 weeks post-conception. We used the cortical surfaces (’.white’ files) provided by the consortium. The segmentation and surface extraction tools are detailed in [Makropoulos and Robinson,]. The quality of the surfaces and corresponding segmentation were checked before the release to confirm the absence of obvious errors.
2) The cortical surface atlas was released with the corresponding article [Bozek et al.,] and is available online  [^2] https://brain-development.org/brain-atlases/atlases-from-the-dhcp-project/cortical-surface-atlas-bozek/. This atlas consists of age-dependant surface meshes spanning 28-44 weeks of gestation, constructed using adaptive kernel regression applied to surface data from the dHCP dataset (described above).
3) The KKI (Kennedy Krieger Institute, [Landman et al.,]) dataset consists of high quality MRI data from 21 healthy adults. The data collection and quality assessment has been described in [Landman et al.,]. We downloaded the raw MRI data and used the recon-all pipeline from freesurfer 6.0 to extract the cortical surfaces as described in [Dale et al.,]. We checked visually the extracted surface for all the 21 subjects to avoid any potential influence of inaccuracy in the surface extraction process.
4) The Human Connectome Project (HCP, [Van Essen and Ugurbil,], S1200 release) dataset consists of the high quality data from 1113 healthy young adults. Structural images were processed using the HCP structural processing pipeline, which has been described in [Glasser and Sotiropoulos,]. We used the preprocessed data (tissue segmentation, cortical surface extraction and registration) released by the consortium.

### v2

# Scale-invariant brain morphometry: application to sulcal depth

@@ Abstract
The geometry of the human cortex is complex and highly variable, with interactions between brain size, cortical folding, and age well-documented in the literature. However, few studies have explored how global brain size influences morphometry features of the cortical surface derived from anatomical MRI. In this work, we focus on sulcal depth, an imaging phenotype that has gained attention in both basic research and clinical applications. We make key contributions to the field by: 1) providing the first quantitative analysis of the influence of brain size on sulcal depth measurements; 2) introducing a novel, scale-invariant method for sulcal depth estimation based on an original formalization of the problem; 3) presenting a validation framework and sharing our code and benchmark data with the community; and 4) demonstrating the biological relevance of our new sulcal depth measure using a large sample of 1,987 subjects spanning the developmental period from 26 weeks post-conception to adulthood.
keywords:
Magnetic resonance Imaging, , cortical surface, , sulcal depth
[…]
organization=Institut de Neurosciences de la Timone, UMR 7289, CNRS, Aix-Marseille Université,addressline=, city=Marseille, postcode=13005, state=, country=France
{graphicalabstract}[Uncaptioned image]
{highlights}
The first quantitative analysis of how brain size affects sulcal depth measurements.
A novel method method based on an original formalization of the problem.
A validation framework with code and benchmark data released.
Experiments on a large sample of 1,987 subjects spanning the developmental period.
@@ 1 Introduction
Sulcal depth is an imaging phenotype that has attracted increasing attention across a wide range of studies, from fundamental research to clinical applications. Key applications include inter-individual cortical surface registration and atlas construction dale_cortical_1999; lyttelton_unbiased_2007; robinson_multimodal_2018, detection of abnormal brain morphology related to clinical conditions auzias_atypical_2014; dierker_analysis_2013, and exploration of the anatomo-functional relationships weiner_mid-fusiform_2013. Furthermore, sulcal depth plays a crucial role in studying individual variations in cortical folding at various spatial scales, from sulci duchesnay_classification_2007 to sulcal basins auzias_deep_2015; im_brain_2008, and even more specific geometrical structures such as plis-de-passages bodin_plis_2021; mangin_plis_2019 or gyral peaks zhang_gyral_2023. Despite its recognized importance for characterizing brain morphology, the sulcal depth is still underutilized compared to other widely used morphological descriptors, such as the gyrification index schaer_surface-based_2008; rabiei_local_2017, curvature-related characteristics pienaar_methodology_2008, and sulcus-based morphometry mangin_-vivo_2010. This limited use can be largely attributed to the inconsistent definitions of sulcal depth used in different studies, which complicates a unified understanding of the literature. Moreover, current methods for estimating sulcal depth often overlook the potential influence of variations in global brain size.
@@ 1.1 On the magnitude of variations in global brain size
The inter-individual variations in global brain size has been extensively described in the literature. The recent large-scale study bethlehem_brain_2022 aggregated morphometry measures extracted from a very large population spanning the entire life, using image processing pipelines such as Freesurfer dale_cortical_1999, or the dHCP pipeline makropoulos_developing_2018. The authors applied advanced non-linear regression techniques to estimate the influence of the age on these morphometry measures. From these curves representative of a large population, we can observe a  4-fold increase in total brain volume between birth and adulthood. In a recent work investigating brain maturation before birth using MRI acquired in utero, studholme_motion_2020 reported a  12-fold increase between the $18^{th}$ week of gestation and birth. So regarding the human brain, variations related to brain maturation can reach an order of magnitude of  48-fold. Regarding inter-individual variations, we can also observe from bethlehem_brain_2022 a  3-fold variation in total brain volume ($[0.5,1.5]$ L) across healthy adults brains. A large body of publication investigated the putative sources of the observed variations across individuals, such as the effect of the sex diaz-caneja_sex_2021; fish_influences_2017, the relationship between depth and cortical thickness and curvature demirci_systematic_2023, or the developmental trajectory duan_individual_2020. On the side of cross-species comparisons, studies such as Van_essen_cerebral_2019 reported  17-fold variation in total brain volume between mature macaques and mature humans. See also demirci_cortical_2022 for an analysis of the variations across 12 primate species covering a wide range of sizes and forms. Taking into account the variations in global brain size when comparing measures across individuals or populations is critical in most applications, either basic or clinically oriented.
@@ 1.2 Features estimated locally are affected by global brain size
It is also crucial to keep in mind that the relationship between global brain size and local measures of brain shape is not restricted to a linear interaction. Studies such as reardon_normative_2018 characterized these complex interactions and reported regionally specific patterns of non linear areal scaling in the human cortex across healthy adults. Such interactions between size and shape (often termed allometry) are very common in biological systems goriely_mathematics_2017. More specifically, systematic interactions between brain size, cortical folding and age have been extensively reported in the literature, both before and after birth rodriguez-carranza_framework_2008; yun_temporal_2020; garcia_dynamic_2018; xu_spatiotemporal_2022; bozek_construction_2018. Such interactions require particular attention in the context of developmental diseases that affect both brain size and cortical folding kapellou_abnormal_2006; germanaud_simplified_2014
@@ 1.3 State-of-the-art of sulcal depth estimation techniques
The most intuitive strategy consists in defining the sulcal depth at each point of a cortical surface by computing the distance between the spatial location of the point of interest and the closest point on the ’convex hull’ of the brain. Already with this intuitive definition, two main ambiguities emerge when moving from the concept to its implementation. First, computing the ’convex hull’ that implicitly corresponds to the 0-level (or reference-level) for the depth measure is not trivial. Indeed, a mathematically correct definition of the convex hull of the cortical surface would not follow closely most concave regions such as the orbital frontal region and the internal temporal region, resulting in a local over estimation of the sulcal depth compared to e.g. the external, more convex regions. The available implementations thus rely instead on a proxy computed by applying a morphological closing of the volumetric segmentation mask of the cortex, which is efficient but question the potential influence between the size of the closing operator and the global size of the brain of interest. Second, there is no unique way to compute the distance to the convex hull and several approaches have been proposed, such as the Euclidean distance in 3D van_essen_surface-based_2004, the geodesic distance along the cortical surface cachia_generic_2003, or an adaptive distance in 3D that follows the geometry of the surface yun_automated_2013. We refer the readers interested in a detailed comparison across these types of approaches to yun_automated_2013 which provides both qualitative and quantitative comparisons.
A second strategy for estimating sulcal depth without considering the convex hull has been introduced in dale_cortical_1999. We refer to this method as SULC in the following. SULC is computed iteratively through an inflation process deforming an initial surface into a smooth, inflated one. Formally, the SULC estimation is obtained as:
$SULC(i)=\int_{0}^{\infty}S_{i}(t)\cdot N_{i}(t)dt$ (1)
@@
A third strategy for estimating sulcal depth is based on two hypotheses relating sulcal depth and mean curvature: (1) the location of the extrema of mean curvature are a first approximation of the location of the extrema of sulcal depth and (2) the frequency ad amplitude of curvature and sulcal depth are closely related (luders_curvature-based_2006, chung_deformation-based_2003).
Finally, a specific method denoted as Depth Potential Function (DPF) has been introduced in boucher_depth_2009 as a combination of the second and third strategies. The DPF has been defined as the solution $D$ of the following regularised screened Poisson equation:
@@
In addition to the specific limitations mentioned above, all the methods from the literature have been designed on adult brains only. To the best of our knowledge, the potential influence of global brain size on the sulcal depth estimation obtained at each vertex of the cortical surface was not considered in the publication introducing these methods. The interactions between global brain size and cortical folding magnitude has been explored rarely, with a few notable exceptions: The excellent work pienaar_methodology_2008, introduces the problem of normalization with respect to global brain size but focuses on curvature and not sulcal depth. In meng_spatial_2014, the authors investigate the evolution of locally deepest points in sulci (sulcal pits) during the first 2 years of life. In order to compensate for the variations in brain size across different subjects at different ages that affect the sulcal depth estimates, the authors propose to adapt the parameters of the sulcal pits extraction algorithm as follows: they perform linear regressions in order to approximate relationship between the parameters of the algorithm and respectively the total area maximum and maximal sulcal depth of the cortical surface. The evaluation of the effectiveness of this approach relies on the observed increase in the number of sulcal pits with age that is consistent across major sulci.This study illustrates the need for better formalization of the impact of global brain size on the depth estimation itself. Indeed, in application studies, global brain size is usually considered as a confounding factor to be controlled as much as possible. For instance, in natu_sulcal_2021 the authors investigate the relationship between functional activations and sulcal depth (using SULC) in the medial ventral temporal sulcus, in human children and adults, and macaques. The potential influence of global brain size is ruled-out by normalizing the sulcal depth estimates according to the deepest point within the sulcus of interest. There is thus a clear need to explicitly investigate more specifically the problem of the intricate variations in global brain size and folding.
1.4 Contributions
In the present work, we contribute to this field of research by 1) providing the first quantitative characterization of the influence of brain size on sulcal depth estimations; 2) introducing a new scale-invariant sulcal depth estimation method based on an original formalization of the problem; 3) proposing a validation framework and sharing with the community the benchmark data; 4) demonstrating the biological significance of our new sulcal depth measure using a large sample of 1,987 subjects, spanning the developmental period from 26 weeks post-conception to adulthood. The source code for this work is available at https://github.com/maximedieudonne/DPF-star, and the benchmark data is available at https://doi.org/10.5281/zenodo.14265228. This article extends our preliminary work dieudonne_scale-controlled_2024.
@@ 2.1 Scale controlled family of functions on surfaces
Note that the scale invariance from Definition 1 is not very relevant in applications to real brains. Indeed, as stated in the section 1, variations between any two brains in global brain size are systematically associated with variations in folding magnitude. When considering two cortical surfaces from different individuals, variations in brain size cannot be approximated by such simple scaling since brain volume and surface are related by an allometry. Even when considering the cortical surfaces from a single individual acquired at two different developmental stages, brain maturation impacts simultaneously both the scaling and the folding. As a consequence, in the vast majority of applications, the experimenter aims at comparing the functions computed on $M_{1}$ and $M_{2}$, knowing that $M_{2}\neq sM_{1}$.
@@ 2.3 Adaptation of the scale of interest through $\alpha$
Now let us focus on the setting of the only one parameter of DPF (Eq.3), and thus of $DPF^{*}$ : $\alpha$. As mentioned in boucher_depth_2009, decomposing the DPF into the eigenfunctions basis of Laplace-Beltrami operator (Fourier modes) allows to better interpret the influence of $\alpha$. The eigenpairs $(\lambda_{i},\Phi_{i})$ are solutions of
@@
This formal analysis allows us to interpret the $DPF^{\star}$ as a low-pass filter on the curvature, the level of which is controlled by $\alpha$. Decreasing $\alpha$ increases the influence of the low frequency components of the curvature compared to high frequency components, and vice versa. The low frequency components correspond to geometric information relative to the global shape of the brain, with large concave regions at the scale of a lobe (such as the prefrontal frontal and occipital regions). Reducing the high frequency components reduces the influence of local folds and noise within the sulci. We refer to germanaud_larger_2012 for further interpretation of $\lambda_{i}$ in terms of wavelengths of an eigenfunction $\Phi_{i}$.
@@ 3 Materials and Experiments
We evaluate our method using quantitative and qualitative comparison with the closest related approach from the literature, the method SULC that was described in Sc.1.3. In this work, we compute SULC using the deform-mesh process available in the bioMedIA mirtk pipeline  [^1] https://github.com/BioMedIA/MIRTK. We confirmed that the estimation is very close to the one from Freesurfer by computing the correlation between the two implementations for several brains and observed a correlation coefficient systematically superior to .99. We present five complementary experiments:
  • 1.
@@
  • 4.
    In the fourth experiment, we compare our method and SULC on the surfacic atlas from bozek_construction_2018 that is representative of the evolution of brain folding with age during normal early development.
  • 5.
    The fifth experiment consists in assessing the robustness of our method with respect to the spatial resolution of the mesh.
@@
1) The third release of the dHCP dataset (http://www.developingconnectome.org edwards_developing_2022) consists of 887 MRI sessions from 783 newborn babies covering the ages from 26 to 45 weeks post-conception. We used the cortical surfaces (’.white’ files) provided by the consortium. The segmentation and surface extraction tools are detailed in makropoulos_developing_2018. The quality of the surfaces and corresponding segmentation were checked before the release to confirm the absence of obvious errors.
2) The cortical surface atlas was released with the corresponding article bozek_construction_2018 and is available online  [^2] https://brain-development.org/brain-atlases/atlases-from-the-dhcp-project/cortical-surface-atlas-bozek/. This atlas consists of age-dependant surface meshes spanning 28-44 weeks of gestation, constructed using adaptive kernel regression applied to surface data from the dHCP dataset (described above).
3) The KKI (Kennedy Krieger Institute, landman_multi-parametric_2011) dataset consists of high quality MRI data from 21 healthy adults. The data collection and quality assessment has been described in landman_multi-parametric_2011. We downloaded the raw MRI data and used the recon-all pipeline from freesurfer 6.0 to extract the cortical surfaces as described in dale_cortical_1999. We checked visually the extracted surface for all the 21 subjects to avoid any potential influence of inaccuracy in the surface extraction process.
4) The Human Connectome Project (HCP, van_essen_human_2012, S1200 release) dataset consists of the high quality data from 1113 healthy young adults. Structural images were processed using the HCP structural processing pipeline, which has been described in glasser_minimal_2013. We used the preprocessed data (tissue segmentation, cortical surface extraction and registration) released by the consortium.